# A Flexible, Equivariant Framework for Subgraph GNNs via Graph Products and Graph Coarsening

### v1

@@ 7 Conclusions
Limitations. Our method operates over a product graph. Although we provide control over the size of this product graph, achieving higher performance requires a larger bag size. This can become a complexity bottleneck, particularly when the original graph is large.
@@ Acknowledgements
The authors are grateful to Beatrice Bevilacqua, for helpful discussions, and constructive conversations about the experiments. HM is the Robert J. Shillman Fellow and is supported by the Israel Science Foundation through a personal grant (ISF 264/23) and an equipment grant (ISF 532/23). FF is funded by the Andrew and Erna Finci Viterbi Post-Doctoral Fellowship.
@@ Appendix A Basic Definitions
Here, $f^{t}$ is an arbitrary (parameterized) continuous function, $\text{agg}^{t}_{i},\ i=1,\dots 4$ are learnable permutation invariant aggregation functions, $e_{v,v^{\prime}},\tilde{e}_{S,S^{\prime}}$ are the (optional) edge features of $G$ and $\mathcal{T}(G)$ respectively and the function $z:\mathcal{P}([n])\times[n]\times\mathcal{P}([n])\times[n]\rightarrow\mathbb{R}^{d}$ maps each tuple of indices $\mathbf{v}=(S,v,S^{\prime},v^{\prime})$ to a vector uniquely encoding the orbit of $\mathbf{v}$ under the action of $S_{n}$ as described in 72.
@@ D.1 Recovering Subgraph GNNs
D.2 Comparison to Theoretical Baseline
@@ Appendix E Linear Invariant (Equivariant) Layer – Extended Section
Solving Equations 65 and 66. Let $\sigma\in S_{n}$ denote a permutation corresponding to the permutation matrix $\mathbf{P}$. Let $\mathbf{P}\star\mathbf{L}$ denote the tensor that results from expressing $\mathbf{L}$ after renumbering the nodes in $V^{\mathcal{T}},V$ according to the permutation $\sigma$. Explicitly, for $\mathbf{L}\in\mathbb{R}^{2^{n}\times n}$, the $(\sigma(S),\sigma(i))$-entry of $\mathbf{P}\star\mathbf{L}$ equals to the $(S,i)$-entry of $\mathbf{L}$. The matrix that corresponds to the operator $\mathbf{P}\star$ in the standard basis, $\mathbf{e}^{(i)}\otimes\mathbf{e}^{(S)}$ is the kronecker product $\mathbf{P}_{\mathcal{I}}\otimes\mathbf{P}_{\mathcal{S}}$. Since $\operatorname{vec}(\mathbf{L})$ is exactly the coordinate vector of the tensor $\mathbf{L}$ in the standard basis we have,
@@
Given Equations 65 and 67 and Equations 66 and 68, it holds that we should focus on solving,
@@ G.1 Proofs of Appendix B
We now restate and prove the proopsitions and Lemmas of Appendix B.
@@
Since we are concerned only with input graphs $G$ from a finite family of graphs (where "finite" means that the maximal graph size is bounded and all possible node and edge feature values come from a finite set), we assume that for any $v\in[n]$, $i\in[k]$, both the input feature vectors $X_{v}\in\mathbb{R}^{d}$ and the encoding vectors $z_{i}\in\mathbb{R}^{d^{*}}$ are one-hot encoded. We aim to show that under these assumptions, any function $f(\cdot)$ of the form 94 can be realized through a single-layer update detailed in Equations 96 , 95, where $M$ is a 4 layer MLP , and $U$ is a single linear layer. The proof involves the following steps:
@@
which encodes each such possible value as a one-hot vector in $\mathbb{R}^{k}$ supported on the last $k-d-d^{*}$ coordinates (this is possible because of equation 106). Now, employing theorem G.1, we define the $x_{j}$s as all possible (distinct) values of 110, with each corresponding $y_{j}$ being the output $\text{enc}(x_{j})$. The theorem now tells us that there exists a 4-layer fully connected ReLU neural network capable of implementing the function $\text{enc}(\cdot)$. We choose $M^{t}$ to be this network. Now since $m^{t}_{i}$, defined in equation 107 is a sum of one-hot encoded vectors, it effectively counts the number of each possible value in the set 108. This proves step 1.
@@
We note that since the output of $\text{agg}^{t}_{i}$ is always within the set $\{0,1\}^{k_{i}}$, the outputs of $V^{t}_{i}$ is always within $\{0,1\}^{k_{1}+\dots+k_{4}}$. Now for $i=1,\dots 4$, employing theorem G.1 we define a dataset $\{x_{j},y_{j}\}_{j=1}^{N}$ by taking the $x_{j}$s as all possible (distinct) values of $\mathcal{X}^{t}(S,v)+m^{t}_{i}$, with each corresponding $y_{j}$ being the output $V^{t}_{i}(x_{j})$. We note that there are finitely many such values as both $\mathcal{X}^{t}(S,v)$ and $m^{t}_{i}$ are one-hot encoded vectors. The theorem now tells us that there exists a a 4-layer fully connected ReLU neural network capable of implementing the function $V^{t}_{i}(\cdot)$. We choose $U^{t}_{i}$ to be this network. Equations 120 - 123 now give us:
@@ G.2 Proofs of Appendix C
To achieve this, we use the layer update as per Definition B.2, which was demonstrated in Proposition B.1 to be equivalent to the general equivariant message passing update in Definition A.5. First, we observe that the graphs $G$ and $H$ are WL-indistinguishable. We then observe that since $|V^{\mathcal{T}}|=1$, the graphs induced by the adjacency matrices $A_{G}$ and $A_{H}$ in Definition B.1 are isomorphic to the original graphs $G$ and $H$, respectively, and therefore they are also WL-indistinguishable. Additionally, we notice that the graphs induced by the adjacency matrices $A_{\mathcal{T}(G)}$ and $A_{\mathcal{T}(H)}$ in Definition B.1 are both isomorphic to the fully disconnected graph with 8 nodes, making them WL-indistinguishable as well. WWe also observe that there exists a bijection $\sigma:V_{G}\rightarrow V_{H}$ that maps all nodes of degree 3 in $G$ to all nodes of degree 3 in $H$. The definition of $\mathcal{T}(\cdot)$ implies that $\sigma$ is an isomorphism between the adjacency matrices $A_{P_{i}}$ corresponding to $G$ and $H$, where $i=1,2$. Finally, we notice that for both $G$, and $H$, the node feature map induced by $\pi_{\text{S}}$ satisfies:
@@
Here the operation $\text{agg}\vline_{a:b}$ initially projects all vectors in the input multi-set onto coordinates $a$ through $b$, and subsequently passes them to the function agg. equations 189 , 190 guarantee that:
@@
equations 202, 204 and 209 now guarantee that:

### v4

@@ 6 Conclusions
Limitations. Our method operates over a product graph. Although we provide control over the size of this product graph, achieving better performance requires a larger bag size. This can become a complexity bottleneck, particularly when the original graph is large.
@@ Acknowledgements
The authors are grateful to Beatrice Bevilacqua, for helpful discussions, and constructive conversations about the experiments. HM is the Robert J. Shillman Fellow and is supported by the Israel Science Foundation through a personal grant (ISF 264/23) and an equipment grant (ISF 532/23). FF is funded by the Andrew and Erna Finci Viterbi Post-Doctoral Fellowship; FF partially performed this work while visiting the Machine Learning Research Unit at TU Wien led by Prof. Thomas Gärtner.
@@ Appendix A Basic Definitions
Here, $f^{t}$ is an arbitrary (parameterized) continuous function, $\text{agg}^{t}_{i},\ i=1,\dots 4$ are learnable permutation invariant aggregation functions, $e_{v,v^{\prime}},\tilde{e}_{S,S^{\prime}}$ are the (optional) edge features of $G$ and $\mathcal{T}(G)$ respectively and the function $z:\mathcal{P}([n])\times[n]\times\mathcal{P}([n])\times[n]\rightarrow\mathbb{R}^{d}$ maps each tuple of indices $\mathbf{v}=(S,v,S^{\prime},v^{\prime})$ to a vector uniquely encoding the orbit of $\mathbf{v}$ under the action of $S_{n}$ as described in 73.
@@ D.1 Recovering Subgraph GNNs
D.2 Comparison to Natural Baselines
@@
We conclude this section with the following proposition, showing there exists coarsening functions which, when combined with CS-GNN, results in an architecture that is strictly more expressive then node-based subgraph GNNs.
Proposition D.4 (CS-GNN can be strictly more expressive then node-based subgraph GNNs).
Let $\mathcal{T}$ be the coarsening function defined by:
$\mathcal{T}(G)=\{\{v\}\mid v\in V\}\cup E\quad G=(V,E).$ (63)
The following holds:
  1. 1.
    Let $G_{1},G_{2}$ be a pair of graphs such that there exists a node-based subgraph GNN model M where $M(G_{1})\neq M(G_{2})$. There exists a CS-GNNmodel $M^{\prime}$ which uses $\mathcal{T}$ such that $M^{\prime}(G_{1})\neq M^{\prime}(G_{2})$.
  2. 2.
    There exists a pair of graphs $G_{1},G_{2}$ such that for any subgraph GNN model $M$ it holds that $M(G_{1})=M(G_{2})$, but there exists a CS-GNNmodel $M^{\prime}$ which uses $\mathcal{T}$ such that $M^{\prime}(G_{1})\neq M^{\prime}(G_{2})$.
This proposition is proved in Appendix G.
@@ Appendix E Linear Invariant (Equivariant) Layer – Extended Section
Solving Equations 66 and 67. Let $\sigma\in S_{n}$ denote a permutation corresponding to the permutation matrix $\mathbf{P}$. Let $\mathbf{P}\star\mathbf{L}$ denote the tensor that results from expressing $\mathbf{L}$ after renumbering the nodes in $V^{\mathcal{T}},V$ according to the permutation $\sigma$. Explicitly, for $\mathbf{L}\in\mathbb{R}^{2^{n}\times n}$, the $(\sigma(S),\sigma(i))$-entry of $\mathbf{P}\star\mathbf{L}$ equals to the $(S,i)$-entry of $\mathbf{L}$. The matrix that corresponds to the operator $\mathbf{P}\star$ in the standard basis, $\mathbf{e}^{(i)}\otimes\mathbf{e}^{(S)}$ is the kronecker product $\mathbf{P}_{\mathcal{I}}\otimes\mathbf{P}_{\mathcal{S}}$. Since $\operatorname{vec}(\mathbf{L})$ is exactly the coordinate vector of the tensor $\mathbf{L}$ in the standard basis we have,
@@
Given Equations 66 and 68 and Equations 67 and 69, it holds that we should focus on solving,
@@ F.5 Additional Results
F.6 Zinc12k Product Graph Visualization
In this subsection, we visualize the product graph derived from the first graph in the Zinc12k dataset. Specifically, we present the right part of Figure 1, for the case of the real-world graphs in the Zinc12k dataset. We perform this visualization for different cluster sizes, $T\in\{2,3,4,5,8,12\}$, which also define the bag size, hence the notation $T$. The nodes in the product graph, $\mathcal{T}(G)\square G$, are $(S,v)$, where $S$ is the coarsened graph node (again a tuple), and $v$ is the node index (of a node from the original graph). For better clarity, we color the nodes $(S,v)$ with $v\in S$ using different colors, while reserving the gray color exclusively for nodes $(S,v)$ where $v\notin S$. The product graphs are visualized in Figures 7, 7, 7, 10, 10 and 10 below.
Figure 5: $T=2$.
Figure 6: $T=3$.
Figure 7: $T=4$.
Figure 8: $T=5$.
Figure 9: $T=8$.
Figure 10: $T=12$.
@@ G.1 Proofs of Appendix B
We now restate and prove the propositions and lemmas of Appendix B.
@@
Since we are concerned only with input graphs $G$ from a finite family of graphs (where "finite" means that the maximal graph size is bounded and all possible node and edge feature values come from a finite set), we assume that for any $v\in[n]$, $i\in[k]$, both the input feature vectors $X_{v}\in\mathbb{R}^{d}$ and the encoding vectors $z_{i}\in\mathbb{R}^{d^{*}}$ are one-hot encoded. We aim to show that under these assumptions, any function $f(\cdot)$ of the form 95 can be realized through a single-layer update detailed in Equations 97 , 96, where $M$ is a 4 layer MLP , and $U$ is a single linear layer. The proof involves the following steps:
@@
which encodes each such possible value as a one-hot vector in $\mathbb{R}^{k}$ supported on the last $k-d-d^{*}$ coordinates (this is possible because of equation 107). Now, employing theorem G.1, we define the $x_{j}$s as all possible (distinct) values of 111, with each corresponding $y_{j}$ being the output $\text{enc}(x_{j})$. The theorem now tells us that there exists a 4-layer fully connected ReLU neural network capable of implementing the function $\text{enc}(\cdot)$. We choose $M^{t}$ to be this network. Now since $m^{t}_{i}$, defined in equation 108 is a sum of one-hot encoded vectors, it effectively counts the number of each possible value in the set 109. This proves step 1.
@@
We note that since the output of $\text{agg}^{t}_{i}$ is always within the set $\{0,1\}^{k_{i}}$, the outputs of $V^{t}_{i}$ is always within $\{0,1\}^{k_{1}+\dots+k_{4}}$. Now for $i=1,\dots 4$, employing theorem G.1 we define a dataset $\{x_{j},y_{j}\}_{j=1}^{N}$ by taking the $x_{j}$s as all possible (distinct) values of $\mathcal{X}^{t}(S,v)+m^{t}_{i}$, with each corresponding $y_{j}$ being the output $V^{t}_{i}(x_{j})$. We note that there are finitely many such values as both $\mathcal{X}^{t}(S,v)$ and $m^{t}_{i}$ are one-hot encoded vectors. The theorem now tells us that there exists a a 4-layer fully connected ReLU neural network capable of implementing the function $V^{t}_{i}(\cdot)$. We choose $U^{t}_{i}$ to be this network. Equations 121 - 124 now give us:
@@ G.2 Proofs of Appendix C
To achieve this, we use the layer update as per Definition B.2, which was demonstrated in Proposition B.1 to be equivalent to the general equivariant message passing update in Definition A.5. First, we observe that the graphs $G$ and $H$ are WL-indistinguishable. We then observe that since $|V^{\mathcal{T}}|=1$, the graphs induced by the adjacency matrices $A_{G}$ and $A_{H}$ in Definition B.1 are isomorphic to the original graphs $G$ and $H$, respectively, and therefore they are also WL-indistinguishable. Additionally, we notice that the graphs induced by the adjacency matrices $A_{\mathcal{T}(G)}$ and $A_{\mathcal{T}(H)}$ in Definition B.1 are both isomorphic to the fully disconnected graph with 8 nodes, making them WL-indistinguishable as well. We also observe that there exists a bijection $\sigma:V_{G}\rightarrow V_{H}$ that maps all nodes of degree 3 in $G$ to all nodes of degree 3 in $H$. The definition of $\mathcal{T}(\cdot)$ implies that $\sigma$ is an isomorphism between the adjacency matrices $A_{P_{i}}$ corresponding to $G$ and $H$, where $i=1,2$. Finally, we notice that for both $G$, and $H$, the node feature map induced by $\pi_{\text{S}}$ satisfies:
@@
Here the operation $\text{agg}\vline_{a:b}$ initially projects all vectors in the input multi-set onto coordinates $a$ through $b$, and subsequently passes them to the function agg. equations 190 , 191 guarantee that:
@@
equations 203, 205 and 210 now guarantee that:
@@
See D.4
Proof.
First, notice that the super-nodes produced by $\mathcal{T}$ are either of size 1, in which case they correspond to nodes, or they are of size two, in which case they correspond to edges. Since an CS-GNNmodel processes feature maps $\mathcal{X}^{t}(S,v)$ where in the initial layer the sset size of $S$ is encoded in $\mathcal{X}^{t}(S,v)$, we can easily use the CS-GNNupdate in Definition A.5 to ignore all values of $\mathcal{X}^{t}(S,v)$ were $|S|=2$ (This can be done by using $f^{t},\text{agg}^{t}_{1},\dots\text{agg}^{t}_{1}$ in Definition A.5 to zero out these values at each update). This means CS-GNNusing $\mathcal{T}$ is able to simulate an CS-GNNupdate with the identity coarsening function, which was shown in Proposition D.1 to be as expressive as GNN-SSWL+ (Definition D.1) which is a maximally expressive node-based subgraph GNN, thus proving part (1) of the proposition. To prove part (2), notice that using the same reasoning as before, an CS-GNNmodel using $\mathcal{T}$ as a coarsening function cal implement an CS-GNNmodel using the edge coarsening function:
$\mathcal{T}^{\prime}(G)=E\quad G=(V,E).$ (217)
An CS-GNNmodel with the identity coarsening function can be interpreted as a GNN-SSWL+ model. Similarly, an CS-GNNmodel using the edge coarsening function $\mathcal{T}^{\prime}$ generalizes the GNN-SSWL+ framework by extending it from node-based subgraph GNNs to edge-based subgraph GNNs. In fact, the same proof in [39, 12] showing that GNN-SSWL+ is at least as expressive as a DSS subgraph GNN using the node deletion policy (see [4] for a definition of the DSS subgraph GNN), can be used to show that CS-GNNusing the edge coarsening function $\mathcal{T}^{\prime}$ is at least as expressive as a DSS subgraph GNN with an edge deletion policy. The latter model was shown in [4] to be able to separate a pair of 3-WL indistinguishable graphs. In contrast, node-based subgraph GNNs were shown in [12] to not be able to separate any pair of 3-WL indistinguishable graphs. Thus, there exists a pair of graphs which CS-GNNusing $\mathcal{T}$ can separate while node-based subgraph GNNs cant, proving part (2) of the proposition.
G.5 Proofs of Appendix E